OpenStudy (anonymous):

A 120-foot-long rope is cut into 3 pieces. The first piece of rope is twice as long as the second piece of rope. The third piece of rope is three times as long as the second piece of rope. What is the length of the longest piece of rope? SHOW WORK PLEASE :]

OpenStudy (anonymous):

Your 2nd sentence says THE FIRST PIECE OF ROPE IS Your 3rd sentence says THE THIRD PIECE OF ROPE IS It never says The Second piece of rope is so it gets x second piece = x first piece = 2x third piece = 3x x + 2x + 3x = 120 6x = 120 x = 20 put back in second piece = 20 first piece = 2(20) = 40 third piece = 3(20) = 60